### v2

@@ Abstract
This position paper argues that the majority of theory of mind benchmarks are broken because of their inability to directly test how large language models (LLMs) adapt to new partners. This problem stems from the fact that theory of mind benchmarks for LLMs are overwhelmingly inspired by the methods used to test theory of mind in humans and fall victim to a fallacy of attributing human-like qualities to AI agents. We expect that humans will engage in a consistent reasoning process across various questions about a situation, but this is known to not be the case for current LLMs. Most theory of mind benchmarks only measure what we call literal theory of mind: the ability to predict the behavior of others. Measuring this kind of reasoning is very informative in testing the ability of agents with self-consistent reasoning. However, it is important to note the distinction between this and what we actually care about when this self-consistency cannot be taken for granted. We call this functional theory of mind: the ability to adapt to agents in-context following a rational response to predictions about their behavior. We find that top performing open source LLMs may display strong capabilities in literal theory of mind, depending on how they are prompted, but seem to struggle with functional theory of mind – even when partner policies are exceedingly simple. Simply put, strong literal theory of mind performance does not necessarily imply strong functional theory of mind performance. Achieving functional theory of mind, particularly over long interaction horizons with a partner, is a significant challenge deserving a prominent role in any meaningful LLM theory of mind evaluation.
@@ 1 Introduction
Many recent papers have been inspired by how humans are evaluated for theory of mind capabilities when evaluating theory of mind in LLMs (Bubeck et al., 2023; Kosinski, 2023; Ma et al., 2023; Street et al., 2024; Strachan et al., 2024). While this does on the surface seem like a logical course of action, it is important to remind ourselves that AI has a tendency to over optimize for its training objectives in a manner that is quite alien to the way the human brain works. For example, in Bubeck et al. (2023) they tout performance of LLMs on human theory of mind tests, but then in the very next section note a noticeable lack of ”process consistency” in LLM explanations. What this means is that LLMs can come up with compelling explanations for what they do that have very little to do with their actual reasoning process. We must therefore proceed with caution about our conclusions when we evaluate LLMs in a way that does not directly align with what we really care about. When we evaluate humans, we typically focus on what we term in this paper to be literal theory of mind, which is their ability to predict the behavior of other agents. As noted by Ma et al. (2023) this can take the form of various abstractions of this behavior including actions, intentions, beliefs, percepts, desires, knowledge, and emotions. However, with humans we take for granted that this prediction of the behavior of other agents will be consistently applied into their own reasoning process when determining their own behavior. The main insight of our paper is that this process consistency cannot be taken for granted when evaluating LLMs.
@@
To be concrete, our position is that: Current LLM theory of mind benchmarks are broken because they only measure literal theory of mind (Definition 2.1) rather than the functional theory of mind (Definition 2.2) that we actually care about our agents displaying in most cases. In Section 4 we demonstrate simple concrete cases where top open source LLMs do display strong literal theory of mind performance that is illusory in nature because it does not also lead to strong functional theory of mind performance. This serves as a case study to highlight the misleading nature of most LLM theory of mind benchmarks even when we do find that agents achieve impressive performance.
@@
Recently, generative AI, particularly in the form of LLM assistants, has been deployed as a tool for a growing variety of real-world use cases where the LLM must interact with a diverse set of people performing a diverse set of tasks. As typically deployed today, these LLMs are only interacting with users at inference time due to the significant computational cost of continuing to train these large models in the context of individual users. As a result, LLMs must learn to adapt their behavior to users in-context based on recorded interaction histories. Indeed, a similar scenario of growing importance involves LLMs interacting with other AI agents through the agentic framework (Shavit et al., 2023). Our work is inspired by the work of Akata et al. (2023) to leverage canonical repeated games from behavioral game theory as a way to asses the adaptation ability of LLMs across the full spectrum of incentive structures. Unfortunately, we find significant deficiencies in the ability of top open source LLMs to adapt to new scenarios in-context, providing a sobering analysis of the ability of LLMs to reliably adapt without continual training of the model weights themselves.
@@
The position of our paper was inspired by an unexpected result in our early experiments. We considered the incredibly simple scenario of an LLM agent playing the classic game Rock, Paper, Scissors against an agent that always plays the action ”Rock” for 100 consecutive rounds. The optimal course of action is for the LLM to respond with the counter to this action ”Paper” as much as possible. If you were in the LLMs shoes how do you think you would do? It is certainly reasonable to take a few exploratory actions, but there would be little excuse for not winning a high percentage of the 100 rounds. What we found is that a vanilla application of common open source LLMs resulted in a policy that chose each action ”Rock”, ”Paper”, and ”Scissor” roughly evenly. This is interesting because this is actually the famous Nash equilibrium solution for this game. However, this is the solution for optimizing your worst case return across any possible opponent. To act in this way against this particular opponent that always plays ”Rock” for 100 consecutive rounds actually demonstrates a profound lack of theory of mind. But it should be emphasized that this LLM agent does actually display a high-level of literal theory of mind. Indeed, it only takes so many rounds of seeing the other agent take the ”Rock” action before the LLM is able to predict it will keep doing so. The problem is that this prediction has nothing to do with its own chosen course of behavior. This is akin to a chatbot that continues to provide a certain response type to a user even when it would predict this response will lead to a reaction where the user expresses that they are displeased with it. We believe this is an important insight for the AI community to grapple with when measuring theory of mind in agents, and formalize the concept of functional theory of mind to accurately describe the lacking capability.
@@ 2 Two Types of Theory of Mind Between Multi-Agent LLM Policies
The Environment. The interaction process between multiple agents is often formalized as a stochastic game (Shapley, 1953) or rather a Markov game in the fully observable setting (Littman, 1994). For generality in this work, we rather consider the setting of partially observable stochastic games (POSGs) (Kuhn, 1953) that generalize POMDPs (Kaelbling et al., 1998) and MDPs (Puterman, 1994) to the multi-agent and decentralized setting. A POSG can be defined as the tuple . Here is a finite set of agents, is a finite set of global states, and is the set of joint actions across agents . is the state transition function based on the joint actions across agents, and is the joint reward function with reward function for each agent . is the joint set of observations with denoting a finite set of observations for each agent. is the function that produces agent observations based on the state. Finally, is the horizon of interactions before termination.
Agent Interaction. We can now consider environment interaction at each step from the perspective of a given focal agent where we will use to denote set of all other agents. At step agent takes action and the other agents take a joint action yielding a transition from to with probability where . The agents then receives its own observation of this state and reward . In the environments we consider in our paper it is assumed that also contains information about the actions of the other agents at the previous step . However, this need not always be the case. In such a POMDP, each agent’s policy generates its action stochastically based on its own interaction history where . Note that policies being defined in this way subsumes the common case where part of the history is discarded for computational or memory efficiency.
LLM History Representations. As we are interested in evaluating LLMs for this problem it is additionally assumed that this history representation must be encoded in text (i.e. LLM tokens). This implies that there must be a function that we have direct access to converting the observations , actions , and rewards to the form of LLM token representations. Moreover, it is important to note that we are interested in learning generalist policies across tasks with LLMs. As a result, following the protocol of Akata et al. (2023), it can also be useful to include any information about the environment tuple to the agent to promote rapid adaptation to new tasks.
@@ 2.1 Disentangling Literal Theory of Mind from Functional Theory of Mind Evaluation
Definition 2.2 provides us with a functional metric for measuring theory of mind in the presence of other agents parameterized by . However, it is still worth considering when the conclusions from this metric will be much different than conclusions from a literal theory of mind metric following Definition 2.1. To do this we must define a new policy that is directly based on the literal theory of mind model prediction such that behavior is chosen to be a rational maximization of the action-value function. In our experiments we simulate such a policy utilizing ground truth knowledge of the payoff structure and use to denote the regret of such a policy. thus represents the best regret that can be possibly achieved by faithfully following the predictions of our theory of mind model under the assumption that the predictions are correct.
@@ 2.2 Initial Results
We begin by conducting experiments on the Rock, Paper, Scissor domain as discussed in the Introduction. LLAMA-2 had previously been found to achieve strong performance on repeated matrix games (Lorè & Heydari, 2023), so we considered initial experiments with the full LLAMA-2 family (Touvron et al., 2023) as well as the competitive Falcon 40B (Almazrouei et al., 2023) and Mixtral models (Jiang et al., 2024). Our results are provided in Table 1. Our prompting strategy for the LLM policy follows what was established for repeated matrix games by Akata et al. (2023) (see Figure 1). Our prompting strategy for the theory of mind model is detailed in Figure 2. The horizon of interaction is steps against a randomly chosen fixed single action policy i.e. always Rock, always Paper, or always Scissors. 95% confidence intervals of the mean estimate across opponent policies are provided. See Appendix B for further details.
@@ 3 Overview of Existing Benchmarks
Frameworks for Machine Theory of Mind. In the field of multi-agent reinforcement learning, theory of mind is generally interpreted as the ability to directly predict the behavior (i.e. the actions) that another agent will take (Lowe et al., 2017; Rabinowitz et al., 2018). However, there can be a number of different ways to abstractly represent this behavior (Ma et al., 2023). For example, abstraction can be applied in the temporal dimension representing behavior as a composition of hierarchical skills (Sutton et al., 1999; Bacon et al., 2017; Riemer et al., 2018b; 2020; Allen et al., 2020; Abdulhai et al., 2021) or goals (Schaul et al., 2015; Andrychowicz et al., 2017; Yang et al., 2018; Kim et al., 2019a; b). Abstraction can also be applied at the agent level in order to represent the abstract actions of groups (Memarian et al., 2022; Touzel et al., 2024). Moreover, as showcased by the games Hanabi (Bard et al., 2020; Nekoei et al., 2021; Malloy et al., 2021a; Nekoei et al., 2023) and Poker (Ganzfried & Sandholm, 2015; Brown & Sandholm, 2019), theory of mind can be directly related to inferring recursive beliefs about unobserved information or knowledge (Moreno et al., 2021). We consider the efficacy of predictions of all of these representations of behavior as examples of measuring literal theory of mind (Definition 2.1). However, it has been recently recognized in the human-computer interaction (HCI) literature that we must go even further to achieve mutual theory of mind for better collaboration with humans (Wang & Goel, 2022; Zhang et al., 2024). In this framework, the mutual shaping of mental representations as well as the functional goal of improved performance on tasks are emphasized. We are inspired by this aspirational goal in our paper, leading us to define a functional measurement of theory of mind performance (Definition 2.2).
@@
Behavioral Game Theory with LLMs. Our analysis of functional theory of mind in LLMs closely follows work considering LLMs in the context of behavioral game theory. However, this work has not featured contextualization of literal theory of mind performance and is not included as an aspect of popular benchmarks focusing on measuring literal theory of mind. Our paper builds off the work of Akata et al. (2023) who considered LLMs playing the Prisoner’s Dilemma and Battle of Sexes. However, we focus on performance playing with simple policies rather than playing with other LLMs. LLMs will generally tend to generate in-distribution interaction histories when they play themselves, so we also find it more interesting to focus on very simple relatively out of distribution partner policies. For the Battle of Sexes game, Akata et al. (2023) found that GPT-4 only could adapt the canonical human-like alternating policy when using a special form of social cognition prompting that considers its own predictions of the actions of others (similar to ”perspective taking” prompting (Xu et al., 2024)). We also tried this approach in our setting and did not find it to be an adequate fix the issues that LLMs experience. The ability for LLMs to play Rock, Paper, Scissors was previously considered by Fan et al. (2024) and the ability for LLMs to play the Prisoner’s Dilemma was previously considered by Lorè & Heydari (2023), but not in the context of theory of mind.

### v3

@@ Abstract
Our paper argues that the majority of theory of mind benchmarks are broken because of their inability to directly test how large language models (LLMs) adapt to new partners. This problem stems from the fact that theory of mind benchmarks for LLMs are overwhelmingly inspired by the methods used to test theory of mind in humans and fall victim to a fallacy of attributing human-like qualities to AI agents. We expect that humans will engage in a consistent reasoning process across various questions about a situation, but this is known to not be the case for current LLMs. Most theory of mind benchmarks only measure what we call literal theory of mind: the ability to predict the behavior of others. However, this type of metric is only informative when agents exhibit self-consistent reasoning. Thus, we introduce the concept of functional theory of mind: the ability to adapt to agents in-context following a rational response to their behavior. We find that many open source LLMs are capable of displaying strong literal theory of mind capabilities, but seem to struggle with functional theory of mind – even with exceedingly simple partner policies. Simply put, strong literal theory of mind performance does not necessarily imply strong functional theory of mind performance or vice versa. Achieving functional theory of mind, particularly over long interaction horizons with a partner, is a significant challenge deserving a prominent role in any meaningful LLM theory of mind evaluation.
@@ 1 Introduction
Many recent papers that evaluate theory of mind in LLMs have been inspired by how humans are evaluated for theory of mind (Bubeck et al., 2023; Kosinski, 2023; Ma et al., 2023; Street et al., 2024; Strachan et al., 2024). Although this course of action seems logical, it is important to recognize that AI has a tendency to over optimize for its training objectives in a manner that is quite alien to the way the human brain works. For example, Bubeck et al. (2023) tout the performance of LLMs on human theory of mind tests, but they note a noticeable lack of “process consistency” in LLM explanations: LLMs can come up with compelling explanations for what they do that have very little to do with their actual reasoning process. We must therefore proceed with caution about our conclusions when we evaluate LLMs in ways that do not directly align with our objectives. When we evaluate humans, we typically focus on what we call in this paper literal theory of mind, which is their ability to predict the behavior of other agents. As noted by Ma et al. (2023), this concept encompasses various abstractions including actions, intentions, beliefs, percepts, desires, knowledge, and emotions. However, with humans we take for granted that the ability to predict the behavior of other agents will be consistently applied within their own reasoning process when determining their own behavior. The main insight of our paper is that this form of process consistency cannot be taken for granted when evaluating LLMs.
@@
Our position is: Current LLM theory of mind benchmarks are broken because they only measure literal theory of mind (Definition 2.1) rather than functional theory of mind (Definition 2.2), which misses a key aspect of evaluating an agent’s ability to choose appropriate actions based on the behavior of another agent. In Section 4 we demonstrate simple, concrete cases where prominent open source LLMs display strong literal theory of mind performance that is illusory in nature because it does not also lead to strong functional theory of mind performance. Our results serve as a case study to highlight the misleading nature of most LLM theory of mind benchmarks even when we do find that agents achieve impressive performance.
@@
Recently, generative AI, particularly in the form of LLM assistants, has been deployed as a tool for a growing variety of real-world use cases where the LLM must interact with a diverse set of people performing a diverse set of tasks. As typically deployed today, these LLMs are only interacting with users at inference time due to the significant computational cost of continuing to train these large models in the context of individual users. As a result, LLMs must learn to adapt their behavior to users in-context based on recorded interaction histories. Indeed, a similar scenario of growing importance involves LLMs interacting with other AI agents through an agentic framework (Shavit et al., 2023). Our work is inspired by the work of Akata et al. (2023) to leverage canonical repeated games from behavioral game theory as a way to asses the adaptation ability of LLMs across a full spectrum of incentive structures. Unfortunately, we find significant deficiencies in the ability of prominent open source LLMs to adapt to new partners in-context, providing a sobering analysis of the ability of LLMs to reliably adapt without continual training of the model weights themselves.
The term “theory of mind” is used in a variety of different ways across different bodies of literature. We use the term as an idiomatic description of behavior, although it can also be interpreted more literally to mean there is an actual artifact called a “theory of mind” that depicts an explicit description or model of an agent’s mental state. Literal theory of mind can then alternatively be considered as theory of mind prediction, and it is possible to achieve optimal precision without forming an explicit description of an agent’s mental state. For example, a model may be conditioning on a variable that is perfectly correlated with the variables another agent is actually considering. The literature on theory of mind in machine learning is generally unconcerned about this distinction. If this prediction ability is achieved, it is not of practical consequence if there actually is a theory of mind model. Likewise, functional theory of mind can alternatively be considered to be theory of mind reasoning. Our assertion is that functional theory of mind is what the AI research community should be focused on creating and evaluating. Notice that this goal requires the ability to adapt to other agents, but does not necessarily require that the agent has an explicit theory of mind model. If we are describing a capability, it is important to focus on the problem it solves, which must be disentangled from the solution method. To define the capability of theory of mind in such a way that an explicit theory of mind model is needed would only make sense if there were kinds of behaviors expressible with model-based approaches that are not expressible with model-free approaches. However, we know that these algorithm classes have the same expressiveness – at least in the case of reinforcement learning (Sutton & Barto, 2018).
@@
The position of our paper was inspired by an unexpected result in our early experiments. We considered the incredibly simple scenario of an LLM agent playing the classic game Rock, Paper, Scissors against an agent that always plays the action “Rock” for 100 consecutive rounds. The optimal course of action is for the LLM to respond with the counter to this action “Paper” as much as possible. It would be reasonable for an agent to take a few exploratory actions, but there would be little excuse for not winning a high percentage of the 100 rounds. What we found is that a vanilla application of common open source LLMs resulted in a policy that chose each action “Rock,” “Paper,” and “Scissor” roughly evenly. This result is interesting because it is actually the famous Nash equilibrium solution for this game. However, this solution optimizes the worst case return across any possible opponent. To act in this way against an opponent that always plays “Rock” for 100 consecutive rounds actually demonstrates a profound lack of theory of mind. But it should be emphasized that this LLM agent does actually display a high-level of literal theory of mind. Indeed, it only takes so many rounds of seeing the other agent take the “Rock” action before the LLM is able to predict it will keep doing so. The problem is that this prediction has nothing to do with its own chosen course of behavior. We believe this is an important insight for the AI community to grapple with when measuring theory of mind in agents, and we formalize the concept of functional theory of mind to accurately describe this often lacking capability.
@@ 2 Two Types of Theory of Mind Between Multi-Agent LLM Policies
The Environment. The interaction process between multiple agents is often formalized as a stochastic game (Shapley, 1953) or rather a Markov game in the fully observable setting (Littman, 1994). For generality in this work, we consider the setting of partially observable stochastic games (POSGs) (Kuhn, 1953) that generalize POMDPs (Kaelbling et al., 1998) and MDPs (Puterman, 1994) to the multi-agent and decentralized setting. A POSG can be defined as the tuple . Here is a finite set of agents, is a finite set of global states, and is the set of joint actions across agents . is the state transition function based on the joint actions across agents, and is the joint reward function with reward function for each agent . is the joint set of observations with denoting a finite set of observations for each agent. is the function that produces agent observations based on the state. Finally, is the horizon of interactions before termination.
Agent Interaction. We can now consider environment interaction at each step from the perspective of a given focal agent where we will use to denote set of all other agents. At step agent takes action and the other agents take a joint action yielding a transition from to with probability where . The agent then receives its own observation of this state and reward . In the environments we consider in this paper it is assumed that also contains information about the actions of the other agents at the previous step . However, this need not always be the case. In such a POMDP, each agent’s policy generates its action stochastically based on its own interaction history where . Note that policies being defined in this way subsumes the common case where part of the history is discarded for computational or memory efficiency.
LLM History Representations. As we are interested in evaluating LLMs for this problem it is additionally assumed that this history representation must be encoded in text (i.e. LLM tokens). This implies that there must be a function that we have direct access to converting the observations , actions , and rewards to the form of LLM token representations. Moreover, it is important to note that we are interested in learning generalist policies across tasks with LLMs. As a result, following the protocol of Akata et al. (2023), it can also be useful to include any known information about the environment tuple to the agent to promote rapid adaptation to new tasks.
@@ 2.1 Disentangling Literal Theory of Mind from Functional Theory of Mind Evaluation
Definition 2.2111The regret can also be considered a form of loss for consistency with Definition 2.1. We chose this notation, rather than , to stay consistent with the common notation for each in the literature. provides us with a functional metric for measuring theory of mind in the presence of other agents parameterized by . However, it is still worth considering when the conclusions from this metric will be much different than conclusions from a literal theory of mind metric following Definition 2.1. To do this we must define a new policy that is directly based on the literal theory of mind model prediction such that behavior is chosen to be a rational maximization of the action-value function. In our experiments we simulate such a policy utilizing ground truth knowledge of the payoff structure and use to denote the regret of such a policy. thus represents the best regret that can be possibly achieved by faithfully following the predictions of our theory of mind model under the assumption that the predictions are correct.
Interesting Tasks for Theory of Mind. Definition 2.2 also provides us with a natural way to characterize how interesting a task is for showcasing theory of mind. Specifically, the task definition must include distributions over the joint policies of the other agents and start states . We can then consider an expectation over two joint policies drawn from the distribution and define as the optimal response to . The metric would then be . This metric requires that the task itself highlights the need to adapt to the specific policies of the other agents. For example, notice that for Rock, Paper, Scissors against arbitrary one action policies, as described in Section 1, this metric evaluates to a significant expected regret per step when the reward for wins is , for ties, and for losses.
@@ 2.2 Initial Results
We begin by conducting experiments on the Rock, Paper, Scissor domain as discussed in Section 1. Our goal here is to conduct an experiment that is as simple as possible so that we can clearly disentangle literal theory of mind from functional theory of mind without additional conflating factors. LLAMA-2 had previously been found to achieve strong performance on repeated matrix games (Lorè & Heydari, 2023), so we considered initial experiments with the full LLAMA-2 family (Touvron et al., 2023) as well as the competitive Falcon 40B (Almazrouei et al., 2023) and Mixtral models (Jiang et al., 2024). Our results are provided in Table 1. Our prompting strategy for the LLM policy follows what was established for repeated matrix games by Akata et al. (2023) (see Figure 1). Our prompting strategy for the theory of mind model is detailed in Figure 2. The horizon of interaction is steps against a randomly chosen fixed single action policy i.e. always Rock, always Paper, or always Scissors. Note that this interaction horizon is 10x longer than prior work (Akata et al., 2023) as we want to make sure behavior converges. 95% confidence intervals of the mean estimate across opponent policies are provided. See Appendix C for further details.
@@ 3 Overview of Existing Benchmarks
Frameworks for Machine Theory of Mind. In the field of multi-agent reinforcement learning, theory of mind is generally interpreted as the ability to directly predict the behavior (i.e. the actions) that another agent will take (Lowe et al., 2017; Rabinowitz et al., 2018). However, there can be a number of different ways to abstractly represent this behavior (Ma et al., 2023). For example, abstraction can be applied in the temporal dimension representing behavior as a composition of hierarchical skills (Sutton et al., 1999; Bacon et al., 2017; Riemer et al., 2018b, 2020; Allen et al., 2020; Abdulhai et al., 2021) or goals (Schaul et al., 2015; Andrychowicz et al., 2017; Yang et al., 2018; Kim et al., 2019a, b). Abstraction can also be applied at the agent level in order to represent the abstract actions of groups (Memarian et al., 2022; Touzel et al., 2024). Moreover, as showcased by the games Hanabi (Bard et al., 2020; Nekoei et al., 2021; Malloy et al., 2021a; Nekoei et al., 2023) and Poker (Ganzfried & Sandholm, 2015; Brown & Sandholm, 2019), theory of mind can be directly related to inferring recursive beliefs about unobserved information or knowledge (Moreno et al., 2021). We consider the efficacy of predictions of all of these representations of behavior as examples of measuring literal theory of mind (Definition 2.1). However, it has been recently recognized in the human-computer interaction (HCI) literature that we must go even further to achieve mutual theory of mind for better collaboration with humans (Wang & Goel, 2022; Zhang et al., 2024). In this framework, the mutual shaping of mental representations as well as the functional goal of improved performance on tasks are emphasized. We are inspired by this aspirational goal in our paper, leading us to define a functional measurement of theory of mind performance (Definition 2.2).
@@
Behavioral Game Theory with LLMs. Our analysis of functional theory of mind in LLMs closely follows work considering LLMs in the context of behavioral game theory. However, this work has not featured contextualization of literal theory of mind performance and is not included as an aspect of popular benchmarks focusing on measuring literal theory of mind. Our paper builds off the work of Akata et al. (2023) who considered LLMs playing the Prisoner’s Dilemma and Battle of Sexes. However, we focus on performance playing with simple policies rather than playing with other LLMs. LLMs will generally tend to generate in-distribution interaction histories when they play themselves, so we also find it more interesting to focus on very simple relatively out of distribution partner policies. For the Battle of Sexes game, Akata et al. (2023) found that GPT-4 only could adapt the canonical human-like alternating policy when using a special form of social cognition prompting that considers its own predictions of the actions of others (similar to “perspective taking” prompting (Xu et al., 2024)). We also tried this approach in our setting and did not find it to be an adequate fix the issues that LLMs experience. The ability for LLMs to play Rock, Paper, Scissors was previously considered by Fan et al. (2024) and the ability for LLMs to play the Prisoner’s Dilemma was previously considered by Lorè & Heydari (2023), but not in the context of theory of mind.
What We are Advocating For. Our main insight is that benchmarks need an interactive component where the theory of mind reasoning performance of LLMs can be assessed separately from predictions about literal theory of mind. Games such as Codenames (Bills et al., 2025), Hanabi (Bard et al., 2020), Taboo (Ashktorab et al., 2020), and Wavelength (Morrison et al., 2025) are good examples of games that would score highly in terms of the relevance metric we proposed in Section 2. In Appendix A, we describe how literal and functional theory of mind could be measured in each of these games. In our experiments, we focus on simple matrix games in which it is even easier to disentangle these concepts and control for conflating factors.